Project

General

Profile

02312014 Minutes

Attending: Gavin, Evan, Nitin, Siva, Tian, Micheal

FD: At 8 am CSS what frozen and cooling was disabled.
Smooth running over the weekend.

7:30 low numi trigger rate plots

Data taking gap - Nearline plots were updating, 10 hour gap, huge data files, with no warnings.

No NDOS monitoring since the multipoint testing (suspect channel map, data format issues).

Neardet - producing empty spill since 3/28 at 3 pm.


Tian 10936 v 3


Nitin - has new ND and NDOS page almost ready. Will send a link around.